The sum of three counting numbers is 24. One of the numbers is half the sum of the
other two, but is twice their difference. What are the numbers?

Answer by richwmiller(17219)

a+b+c=24,
a=1/2(b+c),
a=2*(b-c)
a=8, b=10, c=6